### v1

# Data Management in the Noisy Intermediate-Scale Quantum Era

Abstract.
Quantum computing has emerged as a promising tool for transforming the landscape of computing technology. Recent efforts have applied quantum techniques to classical database challenges, such as query optimization, data integration, index selection, and transaction management. In this paper, we shift focus to a critical yet underexplored area: data management for quantum computing. We are currently in the Noisy Intermediate-Scale Quantum (NISQ) era, where qubits, while promising, are fragile and still limited in scale. After differentiating quantum data from classical data, we outline current and future data management paradigms in the NISQ era and beyond. We address the data management challenges arising from the emerging demands of near-term quantum computing. Our goal is to chart a clear course for future quantum-oriented data management research, establishing it as a cornerstone for the advancement of quantum computing in the NISQ era.
1. Introduction
Data management is crucial in our increasingly data-driven world, exemplified by the widespread use of database systems and the rapid advancement of big data systems (Abadi et al., 2016; Stonebraker and Pavlo, 2024). In recent years, the field of computer science has been energized by the transformative potential of quantum technologies. Quantum computing promises computational capacities far beyond what traditional computers can achieve (Nielsen and Chuang, 2010). However, quantum computing is still in a nascent stage, i.e., the Noisy Intermediate-Scale Quantum (NISQ) era, characterized by quantum computers that are constrained by noise and limited numbers of qubits (Preskill, 2018).
With the ongoing development of quantum computing, new data management challenges naturally emerge. The fundamental differences between quantum and classical computing call for novel data representations to effectively preserve quantum information (Nielsen and Chuang, 2010; Vinkhuijzen et al., 2024). Moreover, many quantum computing tasks are inherently data- and computation-intensive due to the need to handle large-scale quantum states (Boixo et al., 2018), multidimensional quantum data structures (Biamonte, 2020; Young et al., 2023), and error correction codes (Fowler et al., 2012, 2009; Gottesman, 1997; Terhal, 2015). For example, simulating large-scale quantum computation on classical computers involves processing vast amounts of quantum information, leading to significant scalability and optimization challenges (Boixo et al., 2018; Li et al., 2020; Xu et al., 2023; Young et al., 2023). By addressing data management challenges in the NISQ era, the database community has a unique opportunity to significantly enhance the scalability and reliability of quantum technologies, which will advance both research and real-world quantum applications.
These new data management challenges are not yet clearly defined, despite the recent efforts by the database community. Existing work has explored leveraging quantum computers as new hardware to address classical database challenges, such as query optimization (Trummer and Koch, 2016; Fankhauser et al., 2021, 2023; Schönberger, 2022; Schönberger et al., 2023a, 2023, b; Nayak et al., 2023; Winker et al., 2023a), data integration (Fritsch and Scherzinger, 2023), index selection (Kesarwani and Haritsa, 2024; Gruenwald et al., 2023), and transaction management (Bittner and Groppe, 2020; Groppe and Groppe, 2021; Tim Bittner and Sven Groppe, 2020). However, fundamental questions about data management in the NISQ era remain unanswered. Specifically, how should we define and manage data in the context of near-term and future quantum advancements? Given the unique features of quantum computing, such as superposition and entanglement, what are the new considerations to be addressed for effective data management? What data structures and data management systems will best support the development of quantum technologies, particularly given a limited number of qubits, noise, and other challenges of the NISQ era?
While recent vision papers, tutorials, and surveys (Çalikyilmaz et al., 2023; Hai et al., 2024; Yuan et al., 2023, 2024; Winker et al., 2023b) have begun discussing the intersection of data management and quantum computing, they primarily focus on how quantum technologies can accelerate classical database operations. In contrast, our work delves into an equally important yet underexplored area: data management for quantum computing. At the moment, the fundamental concepts, research directions, and problem definitions in this area remain obscure within the database community. This paper initiates the exploration of these critical aspects and lays the foundation for further deeper integration of data management and quantum computing. We will not dive into the basics of quantum computing and information, as ample resources are available (Young et al., 2023; Quist et al., 2024; Xu et al., 2023; Biamonte, 2020). Instead, we focus on outlining the vision for data management paradigms in the NISQ era and identifying potential research challenges that are particularly relevant to the database community, which could have a significant impact on the advancement of today’s quantum technologies.
Contributions. Our contributions are summarized as follows:
  • Fundamentals: We explain quantum data, differentiate it from classical data, and describe the transformation processes between them (Sec. 2).
  • Roadmap: We present our vision for data management research for quantum computing in three paradigms (Sec. 3).
  • Research problems: We elaborate on near-term research questions in data management for quantum computing (Sec. 4).
2. Data in quantum era
Since the 1960s, database management systems (DBMSs) have been developed and operated on classical computers, which rely on bits as fundamental units of data. Classical bits represent binary values of 0 or 1. In the development of quantum computers and their associated software systems, a key challenge is the implementation of qubits (Popkin, 2016), the quantum counterparts of classical bits. Following the principles of quantum mechanics, qubits differ from bits in properties such as superposition and entanglement. These properties fundamentally differentiate quantum data from classical data, which we will explore in this section.
2.1. Classical data vs. quantum data
Classical data is the information that is collected, processed, and stored with traditional computing methods. Today, most of the classical data is stored and queried using database systems such as relational databases, document stores, graph databases, and vector databases (Software, 2024). We refer to quantum data as information collected and processed using quantum computing devices, i.e., computing devices that follow the rules of quantum mechanics to their advantage (Nielsen and Chuang, 2010).  Quantum data is represented by qubits. Next, we list key differences between quantum and classical data below to help understand the unique features of quantum data.
1. Quantum data is probabilistic. Unlike a classical bit, which is 0 or 1, a quantum bit can be in superposition. Mathematically, a zero state and a one state may be represented by a unit vector in the standard basis. That is, the zero state |$\ket{0}$ is represented by the vector $\begin{bmatrix}1\\ 0\end{bmatrix}$ and the one state |$\ket{1}$ is represented by the vector $\begin{bmatrix}0\\ 1\end{bmatrix}$.  A single qubit state, denoted |$\ket{\psi}$, may be represented by a superposition, i.e., a linear combination of |$\ket{0}$ and |$\ket{1}$: |$\ket{\psi}=\alpha\ket{0}+\beta\ket{1}$, for some pair of complex numbers $\alpha,\beta$ called amplitudes, which satisfy $|\alpha|^{2}+|\beta|^{2}=1$.  The probabilistic nature arises when a measurement is performed. When a measurement is performed on the state |$\ket{\psi}$, an outcome 0 is obtained with probability $|\alpha|^{2}$ and 1 obtained with probability $|\beta|^{2}$, and the state permanently changes to the obtained outcome.
Example 2.1 (Superposition).
Consider the qubit in the following superposition:
|$\vspace{-0.2cm}\ket{\psi}=\frac{1}{\sqrt{2}}\ket{0}+\frac{1}{\sqrt{2}}\ket{1}$
When measuring this qubit, there is an equal probability of 50% ($\left|\alpha\right|^{2}=\left|\beta\right|^{2}=\frac{1}{2}$) to get a 0 or 1 as a result.
In classical computing, individual bits can be concatenated to form a bit string, e.g., the three-bit string “010.” Similarly, we may concatenate qubits into a multi-qubit state. In this case, an $n$-qubit state can be represented as a superposition of |$\ket{x}$ for $x\in\{0,1\}^{n}$, or equivalently a vector of $2^{n}$ components.  Later in Sec. 4, we will discuss a three-qubit state in Fig. 2a.
2. Quantum data is fragile. Quantum computers are anticipated to outperform classical computers in solving certain problems. However, with current quantum technology in the NISQ era, quantum resources remain scarce, and quantum data is prone to noise.  Decoherence (Preskill, 1999), is the process where quantum states lose their coherence due to environmental interactions, resulting in the gradual loss of quantum information. Quantum noise, resulting from unintended couplings with the environment, can significantly degrade the performance of quantum computers (Unruh, 1995). Commonly used noise models, such as amplitude damping, phase damping, bit-flip, and phase-flip, mathematically describe how various types of quantum noise lead to decoherence (Nielsen and Chuang, 2010).
3. Quantum data can be entangled. Another difference between qubits and bits is entanglement (Einstein et al., 1935), leading to spooky action at a distance, a phrase coined by Einstein. Imagine two qubits with qubit $A$ being in Amsterdam and qubit $B$ in San Francisco. When qubit $A$ and qubit $B$ are entangled this means their states are correlated. The spooky part is that if we measure qubit $A$ in Amsterdam and find it in a state of 00 (or $1$), we instantaneously know the state of qubit $B$ is 00 (or $1$), even though it is far away in San Francisco. This correlation is maintained independently of the physical distance between them. The next example introduces Bell’s state, a well-known entangled state.
Example 2.2.
(Bell’s state) We extend the one-qubit state in Example 2.1 to a two-qubit state as follows:
|$\ket{\Psi}=\frac{1}{\sqrt{2}}(\ket{00}+\ket{11})=\frac{1}{\sqrt{2}}\begin{bmatrix}1\\ 0\\ 0\\ 1\end{bmatrix}\vspace{-0.1cm}$
2.2. Quantum-classical data transformation
The transformations that map quantum data to classical data, and vice versa, are essential ingredients of quantum computing.  These transformations can be described as quantum algorithms.
Quantum algorithms, circuits, and gates (Nielsen and Chuang, 2010). A quantum algorithm is a carefully designed quantum process that maps input data, classical or quantum, to the desired output data.  The most commonly used model to describe quantum algorithms is the quantum circuit model. A quantum circuit is a sequence of operations, quantum gates, which manipulate qubits to perform computations. Quantum gates are unitary operators. An operator $U$ is unitary if its conjugate transpose $U^{\dagger}$ multiplied by $U$ (and vice versa) equals the identity matrix $I$, denoted as $U^{\dagger}U=UU^{\dagger}=I$. A unitary operator may be composed as a sequence of basic unitary operators. A set of quantum gates is universal if it can approximate any unitary operator to desirable precision.  A widely known universal gate set consists of Pauli gates (including $X,Y,Z$ gates defined by Pauli matrices), the Hadamard gates (see Example 2.3), the $T$ gate, and the controlled-NOT (CNOT) gate.
Example 2.3.
The Hadamard gate (H) is a unitary operator, which maps |$\ket{0}$ to (|$(\ket{0}+\ket{1})/\sqrt{2}$, and |$\ket{1}$ to (|$(\ket{0}-\ket{1})/\sqrt{2}$.  In matrix form, the operator is represented as
$H=\frac{1}{\sqrt{2}}\left[\begin{array}[]{cc}1&1\\ 1&-1\end{array}\right]$
2.2.1. From quantum data to classical data
To solve computational problems using quantum computers, it is essential to extract classical information from quantum states.  In quantum physics, this is done through measurements, which manipulate and convert quantum states (quantum data) into classical outcomes (classical data). For example, measuring the quantum state |$\ket{\psi}$ in Example 2.1 yields a result of 00 with probability $|\alpha|^{2}$ and $1$ with probability $|\beta|^{2}$.
2.2.2. From classical data to quantum data
To load classical data into a quantum computer, the necessary step is encoding the classical data into the states of qubits. Classical data can be encoded as quantum states via encoding methods like basis encoding, amplitude encoding, and many advanced methods (Marin-Sanchez et al., 2023; Weigold et al., 2021a, b, 2022; Nakaji et al., 2022). Basis encoding maps classical bits of 0 and 1 directly onto the quantum states |$\ket{0}$ and |$\ket{1}$. Amplitude encoding encodes classical data into the amplitudes of quantum states. For instance, we have classical data, a normalized vector $x=\begin{bmatrix}a\\ b\end{bmatrix}=\begin{bmatrix}\frac{3}{5}\\ \frac{4}{5}\end{bmatrix}$. We can encode $x$ as a quantum state $|\psi\rangle=a|0\rangle+b|1\rangle=\frac{3}{5}|0\rangle+\frac{4}{5}|1\rangle$.
3. Our Vision: Data Management for Quantum Computing
To facilitate future research on data management for quantum computing, we first sketch the whole landscape in Fig. 1. We categorize this landscape into three distinct paradigms based on how quantum and classical data are transformed and utilized, and the type of hardware involved—whether a quantum or classical computer is employed. Each paradigm not only guides the exploration of quantum research topics leveraging database technologies but also paves the way for new opportunities in database systems, ultimately driving the development of quantum computing.
I Classical paradigm: Classical preprocessing & postprocessing for quantum computing
We begin with the most prevalent setting today: classical computers handle the preprocessing and postprocessing of data for quantum devices, such as quantum chips, quantum sensors, and quantum computers. In this paradigm, data is primarily stored and processed as classical data. For instance, a classical computer may send instructions to a quantum chip. These instructions are classical data and, for example, describe the quantum circuit that should be run. The quantum device runs its computation and, after some time, performs measurements onto one or more of its qubits. The measurement outcomes, which are bits and hence classical data, are then returned to the classical computer. Thus, the quantum device receives classical input (the instructions) and returns the classical output (the measurement outcomes). For some applications, this process is continued as a loop: the measurement output of the current iteration contributes to the next iteration’s input instructions, and both are classical data.
Database research can enhance this process by developing efficient systems for managing, storing, and querying the classical data involved in preprocessing, postprocessing, and iterative feedback loops between classical computers and quantum devices, ensuring seamless integration and optimization of data workflows.
We demonstrate the importance of this paradigm through two key categories of quantum applications: first, applications where the quantum computation is completed immediately upon returning a result. For example, two separated quantum computers can generate a secure, shared key (password) by encoding bits into qubits, then sending and measuring the qubits  (Bennett and Brassard, 1984; Pirandola et al., 2020). This is followed by classical postprocessing on the bitstring (i.e. classical data) that the measurement returns (Elkouss et al., 2011). The resulting bitstring is a secret key that can later be used for secure communication. In the other category, the quantum computation is stalled temporarily, and the quantum chip still holds quantum data on which the computation will continue later. An example is the detection of errors during a quantum computation, where at fixed timesteps during the computation check measurements are performed (Gottesman, 1997; Terhal, 2015), whose outcomes are then decoded to find out which error has most likely occurred (Maurya and Tannu, 2024; Battistel et al., 2023).
Figure 1. Landscape: data management for quantum computing
In this paradigm, often, fast processing of classical data (including I/O with quantum devices) is crucial for practical use, as quantum devices in the NISQ era are noisy, and the quality of quantum data degrades rapidly over time. We delve into specific research challenges, such as error correction using database technologies, in Sec. 4.3.
II Classical-quantum paradigm: classical data represents quantum states and operations.
This hybrid paradigm presents significant potential for new database challenges. It focuses on using classical data to represent and simulate quantum data. For example, Bell’s state from Example 2.2 is represented by a classical vector. Research in this paradigm is more accessible because it does not require a quantum computer; instead, it involves classical computers.
The representative task in this hybrid paradigm is simulation. Simulation is the process of emulating quantum computation, enabling researchers to model and analyze quantum processes as if they were operating on actual quantum hardware (Young et al., 2023; Xu et al., 2023). [^1] In this work, by simulation we refer to classical simulation. Another related term is quantum simulation (Buluta and Nori, 2009; Georgescu et al., 2014), which pertains to simulating quantum mechanics on a quantum computer, a subject studied in physics. Simulations are of paramount importance in the NISQ era (Preskill, 2018). Consider, for instance, the concept of quantum supremacy (Boixo et al., 2018), which seeks to demonstrate the superior capabilities of quantum computers over classical computers. Given that large-scale quantum computers are not yet available in the NISQ era, simulation is essential for comparing the scalability of quantum computers with classical ones. Additionally, simulations play a crucial role in the development of new quantum algorithms, allowing researchers to design, debug, and validate the correctness of these algorithms before deploying them on expensive quantum devices. Moreover, simulations aid in the development of quantum hardware by evaluating error mitigation schemes, predicting algorithm runtimes, and more (Zulehner and Wille, 2018; Jones et al., 2019; Zhou et al., 2020b; Villalonga et al., 2019; Azuma et al., 2021). Simulation serves as a foundational tool across key areas of quantum computing, including quantum supremacy, quantum algorithms, quantum hardware, error correction, and the exploration of potential quantum applications (Xu et al., 2023).
The potential for database research in this paradigm is immense. First, imagine innovative database systems specifically designed to manage classical data that supports simulation. Such systems could handle complex queries, such as updating quantum-state representations as gates are applied, returning classical measurement outcomes, or calculating the closeness of the current quantum state to a desired target state (Vinkhuijzen et al., 2024).
Second, efficient caching strategies for expensive operations, such as repeated simulations of quantum circuits with varying error parameters, present another important research direction. These strategies require the development of efficient data management systems that optimize the computation of such operations.
The third research direction in this paradigm involves the representation of quantum states and operations. Not only can a quantum state be represented as classical data (e.g., vectors), but quantum gates can also be represented as classical data, such as matrices  (van de Wetering, 2020; Hubig et al., 2017). This is useful for quantum design questions such as optimizing the number of quantum gates or gate depth (Kissinger and van de Wetering, 2020; Zulehner et al., 2018) and compiling textbook quantum circuits to only include the low-level operations that real quantum devices support (Smith et al., 2020). As the representation of a quantum state (a vector) or quantum gate (a matrix) generally grows exponentially with the number of qubits, developing efficient methods to store and process these data structures is key to making simulations of quantum computation feasible, either in vector form or through more compressed structures like tensor networks (Orús, 2019) or other advanced representations (Bravyi and Gosset, 2016; van de Wetering, 2020; Zulehner and Wille, 2018; Jozsa and Miyake, 2008).
As we delve deeper into this paradigm, we explore specific challenges and database research opportunities in Sec. 4.1 and 4.2, where we outline the paths forward for leveraging database techniques to meet the emerging demands of quantum computing.
III Quantum paradigm: quantum-native data storage and processing.
Finally, we envision a future research paradigm beyond NISQ, when we have large-scale, fault-tolerant quantum computers. We will deal with pure quantum data, i.e., qubits. Here, the quantum hardware supports storing qubits for long enough to perform other tasks in the meantime. This enables for example a cloud quantum computer (Broadbent et al., 2009) who rotates its resources among multiple clients, various types of quantum machine learning (Cerezo et al., 2022), and Quantum Random Access Memory (Giovannetti et al., 2008), where data can be queried in superposition. A recent vision (Huang et al., 2022) shows a possible future: quantum data is collected from quantum sensing systems, e.g., for discovering a black hole; then stored and processed via the quantum memory of a quantum computer.
The data management research in this paradigm will be centered around handling quantum data, possibly developing quantum-native algorithms. Moreover, storing quantum data will remain challenging, as quantum data storage will still be costly in various ways (the number of qubits is not unbounded, robust storage of quantum data requires large-scale quantum error correction, which is computationally intensive, etc.). Another research topic is to efficiently allocate the available qubits. This includes, for example, various scheduling and allocation tasks (Chandra et al., 2022; Fittipaldi et al., 2023; Meter et al., 2009; Skrzypczyk and Wehner, 2021; Gauthier et al., 2023; Higgott and Breuckmann, 2021) and efficient use of different quantum hardware types each of which has speed-decoherence trade-offs (De Leon et al., 2021). Given the amount of quantum data in this paradigm, most research here will call for novel quantum data processing algorithms and systems beyond NISQ era.
4. Near-Term Research Problems
We will now explore data management research questions across the three paradigms. In Sec. 4.1 and 4.2 we focus on a key challenge within the classical-quantum paradigm. In Sec. 4.3 we broaden the scope to include research opportunities across all three paradigms.
Figure 2. (a) The 3-qubit GHZ state $\frac{1}{\sqrt{2}}(\ket{000}+\ket{111})$ (Nielsen and Chuang, 2010) represented in state vector, whose size is $82^{3}=$ 8. (b) GHZ state represented as a tensor network (Pérez-Garcia et al., 2007; Orús, 2014). For better visualization, in the grey box, we write the tensors in vector/matrix form by slicing the last dimension of A, B, and C. (c) Tensor contraction between matrices $D(i\times k)$ and $E(k\times j)$, leading to a new tensor $F(i\times j)$. (d) The Hadamard gate H (Example 2.3) represented in tensor network. During simulation, the gate operation $H$ executed on qubit state $G$ is represented by contracting $G$ and $H$, leading to a new state $G^{\prime}$.
4.1. The challenge of simulating quantum computation
The input for a simulation is a quantum algorithm described as a quantum circuit, [^2] A uniform family of circuits, to be precise, one for each input size. consisting of input qubits upon which quantum gates are applied, ultimately resulting in a probability vector of the output states, i.e., measurement outcomes (see Sec.  2.2). Then, a simulation is a classical computation that computes that output distribution (so-called strong simulation), or samples from it (weak simulation) (Van Den Nes, 2010). We primarily focus on strong simulation.
The major practical bottleneck of (strong) simulation is scalability (Boixo et al., 2018; Li et al., 2020; Xu et al., 2023; Young et al., 2023). When simulating a quantum state of $n$ qubits, the state is typically represented as a vector with a size of $2^{n}$. The size of the state vector grows exponentially as $n$ increases. For instance, an experiment demonstrating quantum supremacy required 2.25 petabytes for 48 qubits, reaching the memory limits of today’s supercomputers (Boixo et al., 2018). To overcome the memory restriction, existing simulation tools have explored approximation (Vidal, 2003; Medvidović and Carleo, 2021; Jónsson et al., 2018), data compression (Wu et al., 2019), parallelization (Huang et al., 2021), and distributed computing (Smelyanskiy et al., 2016). In Fig. 2 we illustrate concepts mentioned in this section.
4.2. Databases to the rescue
Simulation offers significant opportunities for database research, and conversely, database expertise can greatly enhance simulation.
{myframe}
[roundcorner=3pt,innerleftmargin = 5pt] We envision a classical-quantum simulation system (CQSS) with the following capabilities: (i) automatically providing the most efficient simulation of the input circuit by selecting optimal data structures and operations based on available resources and circuit properties; (ii) operating inherently out-of-core to support the simulation of large circuits that exceed main memory capacity; (iii) ensuring consistency to prevent data corruption and enabling recovery in the event of large-scale simulation crashes; and (iv) improving the entire simulation workflow, including parameter tuning, data collection and querying, exploration and visualization.
At its core, a CQSS must, at a minimum, be capable of evaluating quantum circuits. This primarily involves performing linear algebraic operations, often described in terms of tensor networks. [^3] For simplicity of exposition, we do not consider non-tensor-based simulation methods which represent quantum states by their symmetries, rather than by complex-valued vectors (Gottesman, 1998). A tensor is a multidimensional array, and the dimensions along which a tensor extends are its indices. A tensor network is a collection of tensors represented as nodes, with the shared indices contracted. Tensor contraction is the summation of shared indices between tensors and is a generalization of matrix multiplication (Biamonte, 2020), as shown in Fig. 2c. The first question that comes to mind is as follows.
{myframe}
[roundcorner=3pt,innerleftmargin = 5pt] Q1. Should we push the simulation workload to existing DBMSs?
In other words, should we map the simulation workload to SQL queries and store the results in a DBMS? The advantage is that we can benefit from the efficiency and portability of modern database systems. And indeed, initial efforts from the database community (Blacher et al., 2023; Trummer, 2024) have begun to address the out-of-core simulation challenge by mapping quantum states and gates to SQL queries. For instance, in (Blacher et al., 2023), qubit states and gates are represented as tensor networks, with sparse tensors in COO format [^4] https://docs.scipy.org/doc/scipy/reference/generated/scipy.sparse.coo_matrix.html and Einstein summation operations mapped to SQL queries. Additionally, the recent abstract (Trummer, 2024) introduces the idea of supporting dense vectors using SQL UNION ALL and CASE statements. However, despite these early efforts that support basic quantum states and gates up to 18 qubits and a circuit depth of 18 in separate experiments, the core challenges of representing general quantum states and operations in DBMSs and achieving scalability remain unresolved. Furthermore, preliminary findings question whether query optimizers in existing RDBMSs can effectively optimize SQL queries for simulation workloads (Blacher et al., 2023). Therefore, designing intermediate representations (Karanasos et al., 2020; He et al., 2022) and seeking more effective query optimization, potentially beyond existing techniques (Neumann, 2024), could be crucial and require further exploration.
On the other hand, the connection to tensor computations suggests the following question.
{myframe}
[roundcorner=3pt,innerleftmargin = 5pt] Q2. Should we instead leverage tensor-based database technologies for simulation? Indeed, the database community has a substantial body of work on tensor computation, especially with the recent synergies between databases and machine learning (ML) such as learning ML models over relational data (Khamis et al., 2020; Boehm et al., 2023; Makrynioti and Vassalos, 2019; Zhou et al., 2020a; Paganelli et al., 2023). These studies range from high-level representations to the runtime optimization of ML workflows. At the representation level, numerous studies represent data as tensors (Koutsoukos et al., 2021) and aim to integrate relational and tensor operations (Schleich et al., 2016; Khamis et al., 2018; Huang et al., 2023). Meanwhile, system-building endeavors from the database community (Gao et al., 2017; Luo et al., 2017; Nikolic and Olteanu, 2018; Schleich et al., 2019; Boehm et al., 2020; Nakandala et al., 2020; He et al., 2022; Park et al., 2022; Boehm et al., 2020; Schleich et al., 2023) focus on performance efficiency and scalability, employing optimization techniques during compilation (e.g., operator fusion (Boehm et al., 2018, 2023)) and runtime (e.g., parallelization (Boehm et al., 2014)). Similar endeavours exist in the areas of compilers and High Performance Computing (Ragan-Kelley et al., 2013; Kjolstad et al., 2017; Chen et al., 2018; Jia et al., 2019).
Our envisioned classical-quantum simulation system holds the potential to achieve optimal performance. By integrating relevant optimization techniques from the aforementioned existing systems, we can facilitate the automatic optimization of the underlying tensor network. The actual runtime performance will depend on several factors: the operations within the tensor network, the sparsity of the tensors involved, the memory layouts (i.e., how tensors are stored in memory, such as row-major or column-major formats), and the available hardware (Boehm et al., 2023; Schleich et al., 2023).
While both machine learning and quantum applications share fundamental building blocks like matrices, tensors, and tables, we must address challenges unique to quantum computing. We highlight two key gaps that need to be overcome:
4.2.1. Quantum-specific optimizations
Existing simulators for quantum computing are mostly implemented in Python (cuQuantum team, 2024; Vincent et al., 2022), C++ (AI, 2021; Vincent et al., 2022) or Julia (Luo et al., 2020). There are also dedicated quantum programming languages and domain-specific compilation techniques. See the survey (Heim et al., 2020). In addition, tensor network rewriting techniques have been considered in the ZX-calculus (Coecke and Duncan, 2011). For tensor-based simulators, a defining feature is the optimization of contraction order, finding the optimal contraction path (Young et al., 2023; cuQuantum team, 2024). We refer to (Young et al., 2023) for a more in-depth exposition of simulators and their optimizations. We will need to incorporate the simulation-specific optimizations in our CQSS as well, and in such a way that these also operate out-of-core. This may incur interesting algorithmic challenges. The following question now arises.
{myframe}
[roundcorner=3pt,innerleftmargin = 5pt] Q3. Can we build an optimizer for simulations?
Such an optimizer should, given a circuit, determine the most efficient way to simulate it. For example, when a circuit allows for a tractable simulation (e.g., circuits of bounded treewidth), the optimizer should compile it into an efficient simulation. Similarly, based on the sparsity of (intermediate) quantum states, specialized sparse gate computations may be employed instead of the standard dense computations. Particular to the quantum setting is the noisy character of the quantum computation, which is known to impact the efficiency of simulation (Aharonov et al., 2023; Qi et al., 2020; Pan and Zhang, 2021; Huang et al., 2020, 2021). Hence, an optimizer needs to take noise levels into account. Moreover, the optimizer should consider the downstream task – whether it’s computing precise probabilities for strong simulations or sampling measurement outcomes for weak simulations – and select the appropriate simulation strategy (algorithm) accordingly.
As part of the optimizer, we may also want to check whether two quantum states are exactly or approximately equivalent, analogous to checking equivalent conjunctive queries (Chandra and Merlin, 1977; Kolaitis and Vardi, 1998; Cohen et al., 2007). Having algorithms in place for testing equivalence may help to avoid redundant computations during simulation.
4.2.2. Quantum-specific data representations
Simulating quantum circuits requires precise and efficient representations of quantum states and operations, accounting for the complex nature of quantum mechanics. Ensuring accuracy and scalability as the system grows in complexity is crucial. We mentioned tensor networks as a way to represent the simulation, but various other data representations exist (Vinkhuijzen et al., 2024), e.g., based on algebraic decision diagrams, matrix product states, just to name a few.
{myframe}
[roundcorner=3pt, innerleftmargin = 5pt] Q4. What are good data representations – possibly beyond tensors – for supporting simulations?
An operation on a data structure is tractable if it executes in polynomial time with respect to the input size (Darwiche and Marquis, 2002). The choice of data representation can significantly impact whether certain quantum operations (e.g., gates, measurement explained in Sec. 2.2) are tractable (Vinkhuijzen et al., 2024). We emphasize three main requirements for data representation: expressiveness, closure, and succinctness/tractability.
For expressiveness, the data representation must be sufficiently rich to represent quantum states and measurements. Closure refers to the property that the result of quantum operations on the represented states can also be represented within the same data representation. This property ensures compositionality. Finally, it is crucial to minimize the space required to represent a quantum state, as the number of qubits and the level of entanglement can result in exponentially large states. Therefore, we seek data representations that are succinct, minimizing the space needed to store data while still allowing for efficient query processing. It may also be interesting to consider these questions beyond tensor network representations as well (Bravyi and Gosset, 2016; van de Wetering, 2020; Zulehner and Wille, 2018; Jozsa and Miyake, 2008).
By addressing these challenges, we can create a robust and efficient classical-quantum simulation system that integrates the strengths of existing optimization techniques of databases, tensor computation, and simulations, while meeting the specific needs of quantum computing. Due to space limitations, we will not detail the consistency and recovery aspects of the envisaged CQSS; We comment on its scientific data management part in the next section.
4.3. More data management opportunities
Beyond simulation for quantum computing, quantum-related research is broad, e.g., error correction (Shor, 1995; Steane, 1996; Gottesman, 1997), quantum networks (Wehner et al., 2018; Dahlberg et al., 2019). Next, we expand our discussion to uncover more opportunities spanning all three paradigms shown in Fig. 1.
Quantum error correction & Graph analytics. As discussed in Sec. 2.1, qubits in the NISQ era are inherently vulnerable to noise, which can significantly impact the reliability of computation results. Quantum error correction is crucial for ensuring reliable quantum computation. This process consists of two main phases: coding and decoding. In the coding phase, the input quantum information is coded into a quantum error-correcting code such as the surface code (Fowler et al., 2012; Bravyi et al., 2018; Fowler et al., 2009). During the decoding phase, errors are detected and corrected. Typically, errors are modeled as graphs. The error detection and correction algorithms, referred to as decoders, are often built upon graph theory. For instance, in certain promising error-correction schemes (Fowler, 2015; deMarti iOlius et al., 2024; Wu et al., 2022), decoding involves solving a minimum-weight perfect matching (MWPM) problem on a graph where each node represents a check measurement. Exploring the recent effort of graph analytics (Manolescu and Mohanty, 2023; Yan et al., 2005; Wang et al., 2021; McGregor, 2014; Behnezhad et al., 2023) and graph databases (Tian, 2023) such as Neo4j [^5] https://neo4j.com/, TigerGraph [^6] https://www.tigergraph.com/, might yield valuable insights to advance these error correction techniques. A unique challenge in graph data management for error correction is balancing accuracy and efficiency (deMarti iOlius et al., 2024). A more accurate decoder generally requires higher computational complexity and, therefore, operates more slowly. However, the fidelity of a quantum state decays over time, demanding fast error detection and correction by the decoder.
Quantum experiments & Scientific data management. Quantum experiments, like any other scientific experiments, generate data and require data management. For instance, quantum mechanics experiments on supercomputers often apply HDF5 files to store data and XML files for metadata (Cruz et al., 2012). Typical challenges include avoiding I/O bottleneck (efficient I/O parallelization) and handling the sheer volume and heterogeneity of data files. The scientific data management challenge is also domain-dependent, such as those arising from quantum chemistry or condensed matter physics experiments. These challenges may require us to push forward the frontier of existing database topics such as metadata management (Bernstein et al., 2000; Amano et al., 2014) and data formats (Zeng et al., 2023). Moreover, it is also interesting to explore how to build novel data lakes (Nargesian et al., 2019; Hai et al., 2023) or lakehouses (Armbrust et al., 2020, 2021; Jain et al., 2023) for quantum applications. All these challenges also relate to enabling and improving the simulation workflow of our envisioned classical-quantum simulation system, as described before.
Quantum network & Distributed databases. Quantum networks hold the potential to offer advantages and capabilities beyond those of classical internet systems, e.g., secure communication (Broadbent et al., 2009) or distributed computing (Nickerson et al., 2014). Developing quantum network requires simulators of various kinds, such as simulating timing behaviour accurately (Coopmans et al., 2021; Wu et al., 2021) or for supporting the development of application software (Dahlberg and Wehner, 2018). It is interesting to explore whether the quantum network applications and tools can be supported and advanced by distributed databases (Özsu et al., 1999) and cloud-native databases (Dong et al., 2024), while also considering quantum-specific features like entanglement.
5. Conclusion
We are currently at a privileged time in the evolution of data management, closely aligned with the rise of quantum computing. This convergence calls for innovative approaches to data representation, processing, and querying that are compatible with quantum computing. In this work, we pave the path for future quantum data management by the clarification of the unique features of quantum data compared to classical data, and the exploration of three data management paradigms, which reveal a rich field of complex and significant challenges. As we continue to explore these paradigms, it becomes clear that the challenges we face, such as enhancing simulations with database technologies, are just the beginning. There exists a broader spectrum of challenges that remain unexplored, which require sustained and focused efforts from both the data management and quantum computing communities.